A BLOCK OF MASS 2 kg IS PULLED UP ON A SMOOTH INCLINE OF ANGLE 30 DEGREE WITH THE HORIZONTAL  IF THE BLOCK MOVES WITH AN ACCELERATION OF 1 m/s2, FIND THE POWER DELIVERED BY THE PULLING FORCE AT A TIME 4 sec AFTER THE MOTION STARTS

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Since the block is moving with acceleration1m/s^2, from FBD we have

F-mgsin\theta=ma
\implies F=m(a+gsin\theta)=20(1+10sin30^{\circ})=120N
Due to acceleration of1m/s^2, after 4 seconds, the body will attain velocity ofv=at=4m/s
Hence the power delivered byFis\vec{F}.\vec{v}=Fv=120N\times 4m/s=480W